OpenStudy (anonymous):

Solve . Answers: x = {1, 5/3} x = {7/3, 3} x = {-1, 5/3} No Solution

OpenStudy (anonymous):

sorry whats the ques?

OpenStudy (anonymous):

I'm not understanding the format of your question?

OpenStudy (anonymous):

6x-8 over 2=1

OpenStudy (anonymous):

\[\frac{ 6x-8 }{ 2 } = 1\]

OpenStudy (anonymous):

Step 1; cross multiply

OpenStudy (anonymous):

step 2; add 8 to both sides

OpenStudy (anonymous):

step 3; divide both sides with 6

OpenStudy (anonymous):

I apologize for interfering with your response kanwar. (6x-8)/(2)=1 Multiply each term in the equation by 2. (6x-8)/(2)*2=1*2 Simplify the left-hand side of the equation by canceling the common factors. 2(3x-4)=1*2 Multiply 1 by 2 to get 2. 2(3x-4)=2 Divide each term in the equation by 2. (2(3x-4))/(2)=(2)/(2) Simplify the left-hand side of the equation by canceling the common factors. 3x-4=(2)/(2) Simplify the right-hand side of the equation by simplifying each term. 3x-4=1 Since -4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. 3x=4+1 Add 1 to 4 to get 5. 3x=5 Divide each term in the equation by 3. (3x)/(3)=(5)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=(5)/(3)